OpenStudy (anonymous):

the product of i^7 and i^5

jimthompson5910 (jim_thompson5910):

use the rule x^a*x^b = x^(a+b) to get... i^7*i^5 = i^(7+5) i^7*i^5 = i^(12)

jimthompson5910 (jim_thompson5910):

Now if \(\Large i = \sqrt{-1}\), then... \[\Large i = \sqrt{-1}\] \[\Large i^2 = (\sqrt{-1})^2\] \[\Large i^2 = -1\] \[\Large (i^2)^6 = (-1)^6\] \[\Large i^{2*6} = 1\] \[\Large i^{12} = 1\]